# Model Predictive Control and Reinforcement Learning:A Unified Framework Based on Dynamic Programming

### v1

@@ 1 Introduction
We will describe a conceptual framework for approximate DP, RL, and their connections to MPC, which was first presented in the author’s recent books [Ber20] and [Ber22a]. The present paper borrows heavily from these books, the course textbook [Ber23], the overview papers [Ber21a], [Ber22c], as well as recent research by the author and his collaborators. [^1] Special thanks are due to Yuchao Li for extensive helpful interactions relating to many of the topics discussed in this paper. Early discussions on MPC with Moritz Diehl were also greatly appreciated. The suggestions of Manfred Morari and James Rawlings, as well those of the reviewers, were also very much appreciated.
@@ 1.1 An MPC Problem Formulation
In the context of the present paper, however, because $X$ and $U$ can be arbitrary, it is necessary to use a more general line of analysis and a nontraditional definition of stability. In particular, we say that a policy $\mu$ is stable if
@@ 1.3 Rollout with a Stable Policy
The main difficulty with rollout is that computing the required values of $J_{\mu}\big{(}f(x,\mu(x))\big{)}$ on-line may require time consuming simulation. This is an even greater difficulty for the $\ell$-step lookahead version of rollout, where the required number of base policy values increases exponentially with $\ell$. In this case approximate versions of rollout may be used, such as simplified rollout, truncated rollout, and multiagent rollout; see the books [Ber19], [Ber20], [Ber22a], [Ber23], and the subsequent discussion.
@@ 1.4 Off-Line Training and On-line Play
Alternatively, $\tilde{J}$ may be computed off-line with a problem approximation approach, as the optimal or nearly optimal cost function of a simplified optimization problem, which is more convenient for computation (e.g., a linear-quadratic problem approximation, following linearization of nonlinear dynamics of the original problem). Problem simplifications may include exploiting decomposable structure, reducing the size of the state space, neglecting some of the constraints, and ignoring various types of uncertainties.
@@ 1.5 AlphaZero and TD-Gammon
The architecture of AlphaZero is described in Fig. 3. A comparison with Fig. 2 shows that the architectures of AlphaZero and MPC are very similar. They both involve optimization over a truncated rolling horizon with a terminal cost approximation.
The success of the AlphaZero design framework was replicated by other chess programs such as LeelaChess and Stockfish. It is presently believed that the principal contributor to their success is long lookahead (which uses an efficient on-line play algorithm). The deep neural network-based position evaluator, which is trained off-line with an immense amount of chess data to produce a position evaluator and a player that can select moves “instantly” at any given position, has also contributed to success, although likely to a lesser extent.
@@
The principles of the AlphaZero design have much in common with the earlier TD-Gammon programs of Tesauro [Tes94], [Tes95], [TeG96] that play backgammon (a game of substantial computational and strategical complexity, which involves a large state space, as well as stochastic uncertainty due to the rolling of dice); see Fig. 4. TD-Gammon also uses an off-line neural network-trained terminal position evaluator, and importantly, in its 1996 version, it also extends its on-line lookahead by rollout (simulation with the one-step lookahead player that is based on the position evaluator). Tesauro’s programs stimulated much interest in RL in the middle 1990s, and the last of these programs [TeG96] exhibits different and better play than humans. The rollout algorithm, based on Monte-Carlo simulation, has been a principal contributor to this achievement. [^4] Note that AlphaZero and TD-Gammon are trained to select moves assuming that they play against an adversarial opponent. Their design philosophy would be more closely aligned to MPC, if they were to play against a known and fixed opponent, whose moves can be perfectly predicted at any given position.
A striking empirical observation is that while the neural network used in AlphaZero was trained extensively, the player that it obtained is not used directly during on-line play (it is too inaccurate due to approximation errors that are inherent in off-line neural network training). Instead a separate on-line player is used to select moves, based on multistep lookahead minimization, a limited form of rollout, and a terminal position evaluator that was trained using experience with the off-line player (cf. Fig. 3). The on-line player performs a form of policy improvement, which is not degraded by neural network approximations. As a result, it greatly improves the performance of the off-line player.
@@
    The on-line player of AlphaZero plays much better than its extensively trained off-line player. This is due to the beneficial effect of exact policy improvement with long lookahead minimization, which corrects for the inevitable imperfections of the neural network-trained off-line player, and position evaluator/terminal cost approximation.
@@
    The starting point for the Newton step is based on the results of off-line training, and may be enhanced by longer lookahead minimization and on-line rollout.
@@ 2.3 Visualizing Approximation in Value Space
$F(K)x^{2}$ $=\min_{L\in\Re}\min_{u=Lx}\big{\{}qx^{2}+ru^{2}+K(ax+bu)^{2}\big{\}}$ (1)
$=\min_{L\in\Re}\big{\{}q+bL+K(a+bL)^{2}\big{\}}x^{2}$ (2)
$=\min_{L\in\Re}F_{L}(K)x^{2},$ (3)
where the function $F_{L}(K)$ is the Riccati equation operator for the generic linear policy $\mu(x)=Lx$. Figure 10 illustrates the two-step minimization above, and shows how the graph of the Riccati operator $F$ can be obtained as the lower envelope of the linear operators $F_{L}$, as $L$ ranges over the real numbers.
@@
The preceding argument can be extended to $\ell$-step lookahead minimization to show that a similar Newton step interpretation is possible. Indeed in this case, instead of linearizing $F$ at $\tilde{K}$, we linearize at
@@
i.e., at the result of $\ell-1$ successive applications of $F$ starting with $\tilde{K}$. Each application of $F$ corresponds to a value iteration. Thus the effective starting point for the Newton step is $F^{\ell-1}(\tilde{K})$. Figure 12 depicts the case $\ell=2$.
@@ 2.5 Rollout and Policy Iteration
Finally, we note that rollout, like policy iteration, can be applied universally, well beyond the linear-quadratic/MPC context that we have discussed here. In fact, the main idea of rollout algorithms, obtaining an improved policy starting from some other suboptimal policy, has appeared in several DP contexts, including games; see e.g., Abramson [Abr90], and Tesauro and Galperin [TeG96]. [^7] The name “rollout” was coined by Tesauro in the context of rolling dice to play out a game of backgammon. The use of the name “rollout” has gradually expanded beyond its original context; for example the samples collected through trajectory simulation are referred to as “rollouts” by some authors. The adaptation of rollout to discrete deterministic optimization problems and the principal results relating to cost improvement were given in the paper by Bertsekas, Tsitsiklis, and Wu [BTW97], and were also discussed in the neuro-dynamic programming book [BeT96]. Rollout algorithms for stochastic problems were further formalized in the papers by Bertsekas [Ber97], and Bertsekas and Castañon [BeC99]. Extensions to constrained rollout were given in the author’s papers [Ber05a], [Ber05b]. Rollout algorithms were also proposed in untruncated form within the MPC framework; see De Nicolao, Magni, and Scattolini [DMS98], [MaS04], and followup works.
@@
The author’s book [Ber20] focuses on rollout, policy iteration, and its applications in RL, MPC, and discrete optimization, and refers extensively to the journal literature, which includes a large number of computational studies. These studies discuss variants and problem-specific adaptations of rollout algorithms and consistently report favorable computational experience. The size of the cost improvement over the base policy is often impressive, evidently owing to the fast convergence rate of Newton’s method that underlies rollout.
@@ 2.8 The Importance of the First Step of Lookahead
In practical terms, simplifications after the first step of the multistep lookahead can save a lot of on-line computation, which can be fruitfully invested in extending the length of the lookahead. This insight is supported by substantial computational experimentation, starting with the paper by Bertsekas and Castañon [BeC98], which verified the beneficial effect of certainty equivalence (after the first step) as a rollout simplification device for stochastic problems.
@@ 2.10 How Approximation in Value Space Can Fail and What to Do About It
For an example of broad interest, let us assume that $\tilde{J}$ is obtained by training with data an approximation architecture such as a neural network (e.g., as in AlphaZero and TD-Gammon). Let us also focus on the case of one-step lookahead with no truncated rollout. In this case there are three components that determine the approximation error $\tilde{J}-J^{*}$:
@@
    The power of the architecture, which roughly speaking is a measure of the error that would be obtained if infinite data were available and used optimally to obtain $\tilde{J}$.

### v3

@@ 1 Introduction
We will describe a conceptual framework for approximate DP, RL, and their connections to MPC, which was first presented in the author’s recent books [Ber20] and [Ber22a]. The present paper borrows heavily from these books, the course textbook [Ber23], the overview papers [Ber21a], [Ber22c], as well as recent research by the author and his collaborators. [^1] Special thanks are due to Yuchao Li for extensive helpful interactions relating to many of the topics discussed in this paper. Early discussions on MPC with Moritz Diehl were greatly appreciated. The suggestions of Manfred Morari and James Rawlings, as well as those of the reviewers, were also very much appreciated.
@@ 1.1 An MPC Problem Formulation
In the context of the present paper, however, because $X$ and $U$ can be arbitrary sets, it is necessary to use a more general line of analysis and a nontraditional definition of stability. In particular, we say that a policy $\mu$ is stable if
@@ 1.3 Rollout with a Stable Policy
The main difficulty with rollout is that computing the required values of $J_{\mu}\big{(}f(x,\mu(x))\big{)}$ on-line may require time consuming simulation. This is an even greater difficulty for the $\ell$-step lookahead version of rollout, where the required number of base policy values increases exponentially with $\ell$. In this case, approximate versions of rollout may be used, such as simplified rollout, truncated rollout, and multiagent rollout; see the books [Ber19], [Ber20], [Ber22a], [Ber23], and the subsequent discussion.
@@ 1.4 Off-Line Training and On-line Play
Alternatively, $\tilde{J}$ may be computed off-line with a problem approximation approach, as the optimal or nearly optimal cost function of a simplified optimization problem, which is more convenient for computation (e.g., a linear-quadratic problem approximation, following linearization of nonlinear dynamics of the original problem). Problem simplifications may include exploiting decomposable structure, reducing the size of the state space, neglecting some of the constraints, and ignoring various types of uncertainties. [^4] Two successful applications of problem approximation exploiting decomposable structures, where the author was personally involved, are described in the papers [KGB82] and [MLW24]. Another type of problem approximation, involving the use of some type of certainty equivalence, will be discussed in Section 3.
@@ 1.5 AlphaZero and TD-Gammon
The architecture of AlphaZero is described in Fig. 3. A comparison with Fig. 2 shows that the architectures of AlphaZero and MPC are very similar. They both involve optimization over a truncated rolling horizon with a terminal cost approximation.555 Note that AlphaZero is trained to select moves assuming that it plays against an adversarial opponent. Its design philosophy would be more closely aligned to MPC, if it were to play against a known and fixed opponent, whose moves can be perfectly predicted at any given position. In AlphaZero, the cost function approximation takes the form of a position evaluator, which uses a deep neural network, trained off-line with an immense amount of chess data. The neural network training process also yields a player that can select a move “instantly” at any given chess position, and can be used to assist the on-line lookahead process.
The success of the AlphaZero design framework was replicated by other chess programs such as LeelaChess and Stockfish. It is presently believed that the principal contributor to their success is long lookahead, which uses an efficient on-line play algorithm that involves various forms of tree pruning. The off-line trained position evaluator and player have also contributed to success, although likely to a lesser extent.
@@
The principles of the AlphaZero design have much in common with the earlier TD-Gammon programs of Tesauro [Tes94], [Tes95], [TeG96] that play backgammon (a game of substantial computational and strategical complexity, which involves a large state space, as well as stochastic uncertainty due to the rolling of dice); see Fig. 4. TD-Gammon also uses an off-line neural network-trained terminal position evaluator, and importantly, in its 1996 version, it also extends its on-line lookahead by rollout. Tesauro’s programs stimulated much interest in RL in the middle 1990s, and the last of these programs [TeG96] exhibits different and better play than humans. The rollout algorithm, based on Monte-Carlo simulation, has been a principal contributor to this achievement. [^6] The name “rollout” was coined by Tesauro [TeG96] in the context of backgammon. It refers to simulating/“rolling out” and averaging the scores of many backgammon games, starting from the current position and using the one-step lookahead player that is based on the position evaluator.
A striking empirical observation is that while the neural network used in AlphaZero was trained extensively, the player that it obtained off-line is not used directly during on-line play (it is too inaccurate due to approximation errors that are inherent in off-line neural network training). Instead a separate on-line player is used to select moves, based on multistep lookahead minimization, a limited form of rollout, and a terminal position evaluator that was trained using experience with the off-line player (cf. Fig. 3). The on-line player performs a form of policy improvement, which is not degraded by neural network approximations. As a result, it greatly improves the performance of the off-line player.
@@
    The on-line player of AlphaZero plays much better than its extensively trained off-line player. This is due to the beneficial effect of policy improvement with long lookahead minimization, which corrects for the inevitable imperfections of the neural network-trained off-line player, and position evaluator/terminal cost approximation.
@@
    The starting point for the Newton step is based on the results of off-line training, and can be enhanced by longer lookahead minimization and on-line rollout.
@@ 2.3 Visualizing Approximation in Value Space
$F(K)x^{2}$ $=\min_{L\in\Re}\min_{u=Lx}\big{\{}qx^{2}+ru^{2}+K(ax+bu)^{2}\big{\}}$
$=\min_{L\in\Re}\big{\{}q+bL+K(a+bL)^{2}\big{\}}x^{2}$
$=\min_{L\in\Re}F_{L}(K)x^{2},$
where the function $F_{L}(K)$ is the Riccati equation operator for the generic linear policy $\mu(x)=Lx$. Figure 10 illustrates the two-step minimization of the preceding equation, and shows how the graph of the Riccati operator $F$ can be obtained as the lower envelope of the linear operators $F_{L}$, as $L$ ranges over the real numbers.
@@
The preceding argument can be extended to $\ell$-step lookahead minimization to show that a similar Newton step interpretation is possible (Fig. 12 depicts the case $\ell=2$). Indeed in this case, instead of linearizing $F$ at $\tilde{K}$, we linearize at
@@
i.e., at the result of $\ell-1$ successive applications of $F$ starting with $\tilde{K}$. Each application of $F$ corresponds to a value iteration. Thus the effective starting point for the Newton step is $F^{\ell-1}(\tilde{K})$.
@@ 2.5 Rollout and Policy Iteration
It is important to note that rollout, like policy iteration, can be applied universally, well beyond the linear-quadratic/MPC context that we have discussed here. In fact, the main idea of rollout algorithms, obtaining an improved policy starting from some other suboptimal policy, has appeared in several DP contexts, including games; see e.g., Abramson [Abr90], and Tesauro and Galperin [TeG96]. The adaptation of rollout to discrete deterministic optimization problems and the principal results relating to cost improvement were given in the paper by Bertsekas, Tsitsiklis, and Wu [BTW97], and were also discussed in the neuro-dynamic programming book [BeT96]. Rollout algorithms for stochastic problems were further formalized in the papers by Bertsekas [Ber97], and Bertsekas and Castañon [BeC99]. Extensions to constrained rollout were given in the author’s papers [Ber05a], [Ber05b]. Rollout algorithms were also proposed in nontruncated form within the MPC framework; see De Nicolao, Magni, and Scattolini [DMS98], [MaS04], and followup works.
@@
Finally, we note that the author’s books [Ber20], [Ber22a], [Ber23] provide extensive references to the journal literature, which includes a large number of computational studies. These studies discuss variants and problem-specific adaptations of rollout algorithms and consistently report favorable computational experience. The size of the cost improvement over the base policy is often impressive, evidently owing to the fast convergence rate of Newton’s method that underlies rollout.
@@ 2.7 Double Rollout
2.8 Double Newton Step - Rollout on Top of Approximation in Value Space
Given a quadratic cost coefficient $\tilde{K}$ that defines the policy $\mu(x)=L_{\tilde{K}}x$, it is natural and convenient to consider rollout that uses $\mu$ as a base policy. This can be viewed as rollout that is built on top of approximation in value space. We call this algorithm double Newton step, because it consists of two Newton steps: a first step that maps $\tilde{K}x^{2}$ to $J_{\mu}(x)$ and a second step that maps $J_{\mu}(x)$ to the cost function $J_{\tilde{\mu}}(x)$ of the rollout policy $\tilde{\mu}$ that is produced when the base policy is $\mu$; see Fig. 16.
The double Newton step is much more powerful than the algorithm that performs approximation in value space with two-step lookahead starting from $\tilde{K}$. In particular, both algorithms involve multiple steps for solving the Riccati equation starting from $\tilde{K}$. However, the former algorithm amounts to a Newton step followed by a Newton step, while the later algorithm amounts to a value iteration followed by a Newton step (cf. Fig. 12). For this statement to be correct, $\tilde{K}$ should lie within the region of stability. Such $\tilde{K}$ may obtained by using multiple value iterations, as in the case where a multistep lookahead minimization is performed, i.e.  $\ell>1$ (cf. Fig. 12).
Figure 16: Illustration of a double Newton step. Starting from a quadratic cost coefficient $\tilde{K}$ that defines the policy $\mu$, it uses $\mu$ as a base policy to implement a rollout policy $\tilde{\mu}$.
Note that it is also possible to consider variants of rollout on top of approximation in value space, such as truncated and simplified versions. An important example of the truncated version is the TD-Gammon architecture, where the terminal cost function approximation is constructed off-line by using a neural network.
@@ 2.9 The Importance of the First Step of Lookahead
In practical terms, simplifications after the first step of the multistep lookahead can save a lot of on-line computation, which can be fruitfully invested in extending the length of the lookahead. This insight is supported by substantial computational experimentation, starting with the paper by Bertsekas and Castañon [BeC98], which verified the beneficial effect of certainty equivalence (after the first step) as a rollout simplification device for stochastic problems. On the other hand, implementing imprecisely the minimization of the first step can adversely impact the performance of the multistep lookahead policy. This point is often missed in the design of approximate lookahead minimization schemes, such as Monte Carlo Tree Search.
@@ 2.11 How Approximation in Value Space Can Fail and What to Do About It
For an example of broad interest, let us assume that $\tilde{J}$ is obtained by training with data a neural network (e.g., as in AlphaZero and TD-Gammon). Let us also focus on the case of one-step lookahead with no truncated rollout. In this case there are three components that determine the approximation error $\tilde{J}-J^{*}$:
@@
    The power of the neural network architecture, which roughly speaking is a measure of the error that would be obtained if infinite data were available and used optimally to obtain $\tilde{J}$ by training the given neural network.